# A Novel Use of Pseudospectra in Mathematical Biology: Understanding HPA Axis Sensitivity

### v1

@@ 1 Introduction
Mathematical modeling of the HPA axis has varied over the years, often targeting specific diseases or aligning with experimental observations. These models are typically nonlinear ordinary differential equations (ODEs) or delay differential equations (DDEs). Multiple mechanisms, such as time delays, negative feedback loops, nonlinearities, and varying timescales, can induce oscillations; as overviewed in [7]. The “minimal model” [8] incorporates CRH, ACTH, and cortisol and assumes a differentiable negative feedback effect on ACTH and CRH. Under minimal assumptions, this model demonstrates oscillations through a Hopf bifurcation, albeit with unphysical parameters. Another model [9] features similar components but attributes oscillations to Hill-function type nonlinearities, noting the absence of oscillations with physiologically reasonable parameters. Interestingly, perturbations in this model led to two additional fixed points, indicative of “hypercortisolemic” and “hypocortisolemic” depressive groups, suggesting that mechanisms such as delays, rather than nonlinearity, might drive oscillations. This hypothesis is supported by [10], where a model including ACTH, cortisol, and glucocorticoid receptors demonstrated self-sustained oscillations with sufficient delay, despite CRH being modeled as a constant input.
While a normal form adeptly captures the bifurcation structure, it often fails to align well with empirical data. This issue may persist across the models above, regardless of the chosen model’s correctness. The tension arises from the interaction between non-normality and nonlinearity, where existing tools focus either on one aspect or the other but seldom both. For example, studies in [11, 12] have shown that a non-normal linear operator, despite having all stable eigenvalues, can give rise to a wave packet that interacts with the nonlinearity. Such interactions render first-order approximations, such as normal forms that depict bifurcation phenomena, not representative of the entire solution. Moreover, relying exclusively on linear tools, such as spectra and pseudospectra, can fail to capture the transient behaviors of the complete nonlinear system.
We seek a balanced approach by exploring different linearizations that help us comprehensively understand the complete nonlinear system. We explore transient behaviors for nonlinear systems by computing different types of pseudospectra for various linearizations, gaining different insights from each linearization. We consider the following nonlinear DDE model of the HPA axis:
@@
The paper is organized as follows. Section 2 provides a detailed description of the model, including the motivations for the parameters and the process of de-dimensionalizing the model. Section 3, titled “pseudospectra of the model three ways”, presents the pseudospectra calculated from a time-dependent Jacobian, linearization around the limit cycle, and DMD. In the conclusion, we discuss our results and the contexts in which each can be useful for explaining various experiments and observations. Additionally, we explore how pseudospectra might be used for model validation, i.e., determining whether the perturbed model captures experimental results. We also consider the potential of using data-driven methods to capture the spectrum and pseudospectrum.
@@ 3 Pseudospectra of the model three ways
For instance, consider a finite matrix $A$ whose eigenvalues all have magnitude less than $1$. The linear discrete-time dynamical system $x_{k+1}=Ax_{k}$ is stable since $\lim_{k\rightarrow\infty}\|A^{k}\|\rightarrow\infty$. However, the pseudospectral contours may protrude beyond the unit circle, indicating that there could be some transient growth before decay when the system is perturbed. In the case of pseudospectra calculated on a limit cycle, this could imply that the limit cycle takes longer to stabilize after a perturbation. We will discuss this point below and compute so-called Kreiss constants that capture this behavior.
@@
Often, for nonlinear systems, there is a choice regarding how to obtain the linearization or the matrix $A$ in the above definitions. This is more nuanced in the context of delay-differential equations. Furthermore, the $\epsilon$-pseudospectra form distinct sets; therefore, comparing the pseudospectra of one operator to another is also nuanced. In each section, we provide a short preamble explaining what we hope to understand from this linearization, the method used with numerical details. Then, we present the results and their biological ramifications.
@@ 3.1.1 Method:
We are interested in how the eigenvalues and pseudospectra vary with time (for the selection of base points $(x_{0},y_{0})$) and the parameter $h$ to see if they match the experimental findings of more transient growth regarding upward peaks versus downward peaks of cortisol. We consider the spectral abscissa, defined as $\alpha=\alpha(\tau)=\max_{\lambda\in\mathrm{Sp}(\Delta_{\tau})}\mathrm{Re}(\lambda),$ and the distance to instability, defined as $d_{\tau}=\inf\{\epsilon>0:\exists s\in\mathbb{R},\|\Delta_{\tau}(is)^{-1}\|^{-1}=\epsilon\}$ whenever $\alpha(\tau)<0$. The system in Eq. 3 is (asymptotically) stable if and only if $\alpha(\tau))<0$ [21, Chapter 1]. In that case, $d_{\tau}$ tells us the minimum size of perturbation needed to cause the system to become unstable. We also consider the ratio $-\alpha/d_{\tau}$, whenever $\alpha<0$. This index tells us the ratio between the distance of the spectrum to the imaginary axis and the minimum value of $\|\Delta_{\tau}(z)^{-1}\|^{-1}$ on the imaginary axis. With an abuse of terminology, we shall refer to this as a non-normality index. [^3] Since $\Delta_{\tau}(\lambda)$ is nonlinear in the spectral parameter $\lambda$, the phrase non-normality here is used in the sense of large pseudospectral regions enclosing eigenvalues.
@@ 3.1.2 Results:
Figure 3: (a) – (d): Solution trajectories, spectral abscissa, distance to instability, and ratio $-\alpha/d_{\tau}$. (e) – (h): Pseudospectra of the linearized system for different $\tau$. We display pseudospectra by plotting contours of $\epsilon$ (colobars). The eigenvalues are shown in red. The stability line (imaginary axis) is the dotted magenta line.
@@
The index $-\alpha_{\tau}/d_{\tau}$ has peaks before and after the unstable region (panel (d)). These correspond to the dips in the distance to instability (panel (c)) and the spectral abscissa (panel (b)). Therefore, a small perturbation can easily bring us into the unstable half-plane, and this in itself should be considered as a “non-normal feature”. In [23], it was said that the HPA axis needs to be both “sensitive to environmental perturbations, and able to respond differently to both small and large stimuli”, as well as be “robust with preservation of dynamic behaviour during these perturbations”. We can associate the two different sized peaks with different responses to small and large stimuli at different times. The fact that there is a plateau in between these two peaks possibly corresponds to robustness and the system’s desire to return to equilibrium following a perturbation but being sufficiently “non-normal” at other points to respond to perturbation. These ideas have to be experimentally validated.
@@ 3.2 Floquet pseudospectrum: Linearization about limit cycle
We now consider a periodic linear DDE system, corresponding to linearization about the limit cycle. Unlike linear periodic ODE systems, a complete Floquet theory does not generally exist. Nevertheless, we can still characterize stability and transients through spectral properties of an infinite-dimensional operator. [^4] In our case, this operator is completely continuous. This implies that its spectrum is countable, zero is the only accumulation point of the spectrum, and all non-zero parts of the spectrum are eigenvalues of finite multiplicity.
@@ 3.2.1 Method:
If the base trajectory $(x_{0},y_{0})$ is periodic, then Eq. 3 is a linear periodic DDE system, where the matrices $A$, $B$ and $C$ are periodic in time. The stability of such systems is characterized by characteristic multipliers [25, Chapter 8], which are similar to those found in Floquet theory. [^5] It should be noted that a complete Floquet theory does not generally exist. Let $\omega$ denote the period of the limit cycle. Consider the operator $U$, which maps a continuous function $\phi$ defined over the interval $[-t_{1},0]$ to the solution of Eq. 3 over the interval $[-t_{1}+\omega,\omega]$, starting from the initial condition $\phi$. The operator $U$ acts on the Banach space $C([-t_{1},0])$ of continuous functions on $[-t_{1},0]$ equipped with the $L^{\infty}$ norm. The stability and transient behavior of the system are determined by the spectra and pseudospectra of $U$. Pseudospectra of $U$ are given by [^6] A subtlety arises here. The spectrum of $U$ is independent of the initial time for which we study the system in Eq. 3 and the eigenspaces for different initial times are diffeomorphic. However, the pseudospectra can change. For consistency, as we vary $h$, we take a point at which ACTH peaks as the initial time.
@@ 3.3.2 Results:
The physical interpretation of these results aligns with those in Section 3.2, in that it is physically reasonable for the system to be stable while exhibiting large regions of pseudospectra. However, the Kreiss constants vary significantly between the methods and show a different type of dependency on $h$. We should not be alarmed by the differences in the Kreiss constants, as the norms they are based on are fundamentally different and incomparable. Furthermore, the Koopman operator provides a global linearization of a dynamical system, in contrast to the local linearization discussed in Section 3.2.
@@ 4 Conclusion
Our first method, the time-dependent Jacobian, provided physically relevant results consistent with early rat experiments [16, 17], where aggressive behavior was observed when rats were perturbed on the upward slope of corticosterone secretion. Additionally, fluctuations in the non-normality index indicated that the rhythms of ACTH and cortisol could produce a range of responses within a single period. Linearization around the limit cycle showed that, although the cycle was stable, the response to perturbations could be disproportionate, as evidenced by the pseudospectra. The DMD method similarly demonstrated stability across all eigenvalues and limit cycle-like structures in the eigenvectors yet highlighted sensitivity to perturbations in the pseudospectra. For computing pseudospectra in the latter two cases, we developed novel mathematical techniques for Banach spaces and for DMD in delay differential equations. Interestingly, increasing $h$ resulted in diminishing transient effects in local linearization, while enhancing them in global linearization, as captured by the Kreiss constants.
@@
Finally, data-driven approaches like DMD help bypass challenges in model substantiation but introduce difficulties in distinguishing perturbations from underlying rhythms or deriving perturbations from data. With established prior rhythms and modelable perturbations, we can investigate how structured pseudospectra on matrices align models with reality. This opens avenues for studying individual variability in perturbation responses and advancing personalized medicine. We advocate for strong collaboration between mathematicians and experimentalists to ensure that data suits these sophisticated techniques.

### v2

@@ 1 Introduction
Mathematical modeling of the HPA axis has varied over the years, often targeting specific diseases or aligning with experimental observations. These models are typically nonlinear ordinary differential equations (ODEs) or delay differential equations (DDEs). Multiple mechanisms, such as time delays, negative feedback loops, nonlinearities, and varying timescales, can induce oscillations, as overviewed in [7]. The “minimal model” [8] incorporates CRH, ACTH, and cortisol and assumes a differentiable negative feedback effect on ACTH and CRH. Under minimal assumptions, this model demonstrates oscillations through a Hopf bifurcation, albeit with unphysical parameters. Another model [9] features similar components but attributes oscillations to Hill-function type nonlinearities, noting the absence of oscillations with physiologically reasonable parameters. Interestingly, perturbations in this model led to two additional fixed points, indicative of “hypercortisolemic” and “hypocortisolemic” depressive groups, suggesting that mechanisms such as delays, rather than nonlinearity, might drive oscillations. This hypothesis is supported by [10], where a model including ACTH, cortisol, and glucocorticoid receptors demonstrated self-sustained oscillations with sufficient delay, despite CRH being modeled as a constant input.
While a normal form adeptly captures the bifurcation structure, it often fails to align well with empirical data. This issue may persist across the models above, regardless of the chosen model’s correctness. The tension arises from the interaction between non-normality and nonlinearity, where existing tools focus either on one aspect or the other but seldom both. For example, studies in [11, 12] have shown that a non-normal linear operator, despite having stable eigenvalues, can give rise to a wave packet that interacts with the nonlinearity. Such interactions render first-order approximations, such as normal forms that depict bifurcation phenomena, not representative of the entire solution. Moreover, relying exclusively on linear tools, such as spectra and pseudospectra, can fail to capture the transient behaviors of the complete nonlinear system.
We seek a balanced approach by exploring different linearizations that help us comprehensively understand the complete nonlinear system. We explore transient behaviors for nonlinear systems by computing different types of pseudospectra for various linearizations, gaining different insights from each linearization.
We consider the following nonlinear DDE model of the HPA axis:
@@
The paper is organized as follows. Section 2 provides a detailed description of the model, including the motivations for the parameters and the process of de-dimensionalizing the model. Section 3, titled “pseudospectra of the model three ways”, presents the pseudospectra calculated from a time-dependent Jacobian, linearization around the limit cycle, and DMD. In the conclusion, we discuss our results and the contexts in which each can be useful for explaining various experiments and observations. Additionally, we explore how pseudospectra might be used for model validation, i.e., determining whether the perturbed model captures experimental results. We also consider the potential of using data-driven methods to capture the spectrum and pseudospectrum. Code producing all of the figures of the paper can be found here: https://github.com/miinadietrich/hpa˙axis˙pseudospectra.
@@ 3 Pseudospectra of the model three ways
For instance, consider a finite matrix $A$ whose eigenvalues all have magnitude less than $1$. The linear discrete-time dynamical system $x_{k+1}=Ax_{k}$ is stable since $\lim_{k\rightarrow\infty}\|A^{k}\|\rightarrow 0$. However, the pseudospectral contours may protrude beyond the unit circle, indicating that there could be some transient growth before decay when the system is perturbed. In the case of pseudospectra calculated on a limit cycle, this could imply that the limit cycle takes longer to stabilize after a perturbation. We will discuss this point below and compute so-called Kreiss constants that capture this behavior.
@@
Often, for nonlinear systems, there is a choice regarding how to obtain the linearization or the matrix $A$ in the above definitions. This is more nuanced in the context of delay-differential equations. Furthermore, the $\epsilon$-pseudospectra form distinct sets; therefore, comparing the pseudospectra of one operator to another is also nuanced. In each section, we provide a short preamble explaining what we hope to understand from this linearization and the method used with numerical details. Then, we present the results and their biological ramifications.
@@ 3.1.1 Method:
We are interested in how the eigenvalues and pseudospectra vary with time $\tau$ (for the selection of base points $(x_{0},y_{0})$) and the parameter $h$ to see if they match the experimental findings of more transient growth regarding upward peaks versus downward peaks of cortisol. We consider the spectral abscissa, defined as $\alpha=\alpha(\tau)=\max_{\lambda\in\mathrm{Sp}(\Delta_{\tau})}\mathrm{Re}(\lambda),$ and the distance to instability, defined as $d_{\tau}=\inf\{\epsilon>0:\exists s\in\mathbb{R},\|\Delta_{\tau}(is)^{-1}\|^{-1}=\epsilon\}$ whenever $\alpha(\tau)<0$. The system in Eq. 3 is (asymptotically) stable if and only if $\alpha(\tau)<0$ [21, Chapter 1]. In that case, $d_{\tau}$ tells us the minimum size of perturbation needed to cause the system to become unstable. We also consider the ratio $-\alpha/d_{\tau}$, whenever $\alpha<0$. This index tells us the ratio between the distance of the spectrum to the imaginary axis and the minimum value of $\|\Delta_{\tau}(z)^{-1}\|^{-1}$ on the imaginary axis. With an abuse of terminology, we shall refer to this as a non-normality index. [^3] Since $\Delta_{\tau}(\lambda)$ is nonlinear in the spectral parameter $\lambda$, the phrase non-normality here is used in the sense of large pseudospectral regions enclosing eigenvalues.
@@ 3.1.2 Results:
Figure 3: (a) – (d): Solution trajectories, spectral abscissa, distance to instability, and ratio $-\alpha/d_{\tau}$. (e) – (h): Pseudospectra of the linearized system for different $\tau$. We display pseudospectra by plotting contours of $\epsilon$ (colorbars). The eigenvalues are shown in red. The stability line (imaginary axis) is the dotted magenta line.
@@
The index $-\alpha_{\tau}/d_{\tau}$ has peaks before and after the unstable region (panel (d)). These correspond to the dips in the distance to instability (panel (c)) and the spectral abscissa (panel (b)). Therefore, a small perturbation can easily bring us into the unstable half-plane, and this in itself should be considered as a “non-normal feature”. In [23], it was said that the HPA axis needs to be both “sensitive to environmental perturbations, and able to respond differently to both small and large stimuli”, as well as be “robust with preservation of dynamic behavior during these perturbations”. We can associate the two different-sized peaks with different responses to small and large stimuli at different times. The fact that there is a plateau between these two peaks possibly corresponds to robustness and the system’s desire to return to equilibrium following a perturbation but being sufficiently “non-normal” at other points to respond to perturbation. These ideas have to be experimentally validated.
@@ 3.2 Floquet pseudospectrum: Linearization about limit cycle
We now consider a periodic linear DDE system, corresponding to linearization about the limit cycle. Unlike linear periodic ODE systems, a complete Floquet theory does not generally exist. Nevertheless, we can still characterize stability and transients through the spectral properties of an infinite-dimensional operator. [^4] In our case, this operator is completely continuous. This implies that its spectrum is countable, zero is the only accumulation point of the spectrum, and all non-zero parts of the spectrum are eigenvalues of finite multiplicity.
@@ 3.2.1 Method:
If the base trajectory $(x_{0},y_{0})$ is periodic, then Eq. 3 is a linear periodic DDE system, where the matrices $A$, $B$ and $C$ are periodic in time. The stability of such systems is characterized by characteristic multipliers [25, Chapter 8], which are similar to those found in Floquet theory. [^5] It should be noted that a complete Floquet theory does not generally exist. Let $\omega$ denote the period of the limit cycle. Consider the operator $U$, which maps a continuous function $\phi$ defined over the interval $[-t_{1},0]$ to the solution of Eq. 3 over the interval $[-t_{1}+\omega,\omega]$, starting from the initial condition $\phi$. The operator $U$ acts on the Banach space $C([-t_{1},0])$ of continuous functions on $[-t_{1},0]$ equipped with the $L^{\infty}$ norm. The stability and transient behavior of the system are determined by the spectra and pseudospectra of $U$. Pseudospectra of $U$ are given by [^6] A subtlety arises here. The spectrum of $U$ is independent of the initial time for which we study the system in Eq. 3, and the eigenspaces for different initial times are diffeomorphic. However, the pseudospectra can change. For consistency, as we vary $h$, we take a point at which ACTH peaks as the initial time.
@@ 3.3.2 Results:
The physical interpretation of these results aligns with those in Section 3.2 in that it is physically reasonable for the system to be stable while exhibiting large regions of pseudospectra. However, the Kreiss constants vary significantly between the methods and show a different type of dependency on $h$. We should not be alarmed by the differences in the Kreiss constants, as the norms they are based on are fundamentally different and incomparable. Furthermore, the Koopman operator provides a global linearization of a dynamical system, in contrast to the local linearization discussed in Section 3.2.
@@ 4 Conclusion
Our first method, the time-dependent Jacobian, provided physically relevant results consistent with early rat experiments [16, 17], where aggressive behavior was observed when rats were perturbed on the upward slope of corticosterone secretion. Additionally, fluctuations in the non-normality index indicated that the rhythms of ACTH and cortisol could produce a range of responses within a single period. Linearization around the limit cycle showed that, although the cycle was stable, the response to perturbations could be disproportionate, as evidenced by the pseudospectra. The DMD method similarly demonstrated stability across all eigenvalues and limit cycle-like structures in the eigenvectors yet highlighted sensitivity to perturbations in the pseudospectra. For computing pseudospectra in the latter two cases, we developed novel mathematical techniques for Banach spaces and for DMD in delay differential equations. Interestingly, increasing $h$ resulted in diminishing transient effects in local linearization while enhancing them in global linearization, as captured by the Kreiss constants.
@@
Finally, data-driven approaches like DMD help bypass challenges in model substantiation but introduce difficulties in distinguishing perturbations from underlying rhythms or deriving perturbations from data. With established prior rhythms and modelable perturbations, we can investigate how structured pseudospectra on matrices align models with reality. This opens avenues for studying individual variability in perturbation responses and advancing personalized medicine. We advocate for strong collaboration between mathematicians and experimentalists to ensure data suits these sophisticated techniques.